OpenStudy (anonymous):

How many solutions does the following system have? y = -2/3 x + 3 and 2x + 3y = 9

OpenStudy (anonymous):

2 solutions

OpenStudy (anonymous):

It would most likely be one. Give me a moment. I shall give you some simple steps. Make a table with x and y. Then try to see if they are equal to the answer given. Ehuman, did you not just give order to give steps?

OpenStudy (anonymous):

solve for x, solve for y x solution plus y solution = 2 solutions

OpenStudy (anonymous):

I was going to get to that once we started, telling him how many solutions doesn't give him any answers, unless that is the whole of the question.

OpenStudy (anonymous):

@jakerottttte is your question complete? y = -2/3 x + 3 and 2x + 3y = 9 For part two, is it x=2x+3y=9, or is it y=2x+3y=9

OpenStudy (jdoe0001):

or just solve the 2nd one for "y" and compare the slopes

OpenStudy (jdoe0001):

@jakerottttte solve the 2nd one for "y", what do you get for the slope?

OpenStudy (jdoe0001):

or just solve the 2nd one for "y" and compare the slopes